OpenStudy (anonymous):

Need help with graphing this Graph the exponential function y=4 (3)x

OpenStudy (anonymous):

its very easy to graph just put it in a calculator

OpenStudy (anonymous):

I don't have a calculator right now

OpenStudy (anonymous):

online calculator @MsJasmine

OpenStudy (anonymous):

When you're graphic something, just plug in the X into the equation. So for X = 1, you plug it in, and it is 4*3*1 or 12, so at X=1, you go all the way up to Y=12 and plot that point. Then you repeat this until you get all the points.

OpenStudy (anonymous):

where can I find one at do you no of a website

OpenStudy (anonymous):

Type into Google "Graph 4 (3)x"

OpenStudy (anonymous):

i use https://www.desmos.com/calculator it helps me a lot

OpenStudy (anonymous):

ok so when I do it I don't put the y just 4 (3)x

OpenStudy (anonymous):

Yeah, you can also put the Y but Google already detects what you're trying to do either way.

OpenStudy (anonymous):

ok let me try it

OpenStudy (anonymous):

its not showing me how the graph looks its just showing me how to put it in I went to google an put it in the search

OpenStudy (anonymous):

It should look something like this: http://puu.sh/5Mthj.png

OpenStudy (anonymous):

yes but the graphs on my work is not the same they look like this

OpenStudy (anonymous):

Are you missing an exponent in your original question?

OpenStudy (anonymous):

no it says Graph the exponential function y=4 (3)x

OpenStudy (anonymous):

can anyone help me with this one

OpenStudy (anonymous):

If you're sure that's what the graph looks like, then you're writing the equation wrong. Can you click "draw" below the comment box and write out the problem instead?

OpenStudy (anonymous):

Im sure that's how it looks ill post it again... Graph the exponential function y=4 (3)x first one is A next is B then C and D

OpenStudy (anonymous):

I picked d an it was wrong

OpenStudy (anonymous):

can someone help me

OpenStudy (wolf1728):

If the function is exponential it should be a curve (which they are) but the equation can't be y=4 (3)x because that would be just a line. Could you check that equation again?

OpenStudy (anonymous):

 THATS HOW IT LOOKS